A particle of mass m projected vertically with a speed u. If air resistance is R then find out time of ascent, time of descent,height and the speed with which it will return to the point of projection?

Grade:12th pass

1 Answers

Himanshu Rathour
5 years ago
For ascent
Net acceleration=g+R.    (g is downwards and air resistance will oppose the motion)
Now, using v=u+at
t=u/(g+R) 
Height. H=u²/2(g+R)
For descent
Net acceleration will be (g-R)    {g will be downwards and R will be upwards)
Now using h=ut +1/2at²
u²/2(g+R)=1/2(g-R)t²
t=u/(g²-R²)½
Hence
Time of ascent =u/(g+R) 
Time of descent =u/(g²-R²)½
Height=u²/2(g+R)